# Artificial Neural Networks for Magnetoencephalography: A review of an emerging field

### v1

@@ 3.4 Other studies
This third category includes all the studies we found that combine MEG with ANNs but do not technically fall into the classification or modeling categories. Generally speaking, these studies cover a variety of topics such as data preprocessing techniques ([104, 105, 106, 107, 108, 109, 110, 111]), source estimation (solving the MEG inverse problem) ([128, 129, 130, 131, 132, 134, 135, 133, 146]) and pure methods-oriented articles ([113, 114, 115, 116, 117, 147, 124, 119, 118, 120, 127, 121, 123, 126, 125]).
Studies that employ ANN to tackle the MEG inverse problem (i.e. estimating the cortical sources from the sensor level data) typically employ simulated data from known sources, projecting them onto M/EEG sensor space with varying signal-to-noise ratios (SNRs). This simulated data is then used as input for an artificial neural network (ANN) pretrained to predict the original source location. Performance evaluation commonly involves assessing peak or average location error, often in conjunction with Area Under the Curve (AUC) metrics or spatial dispersion measures with AUC. Source estimation papers frequently use simulated data first to train their network, and then further validate their models on actual MEG data and compare their performance to other established techniques such as Minimum Norm Estimate (MNE) ([148]), Beamformer ([149]) or e- or s-LORETA ([150]). Studies in this subcategory exclusively use RNNs, ResNets, or CNNs, sometimes augmented with attention. Overall, the application of ANNs to source localization shows promise. Recently introduced novel baseline methods like ConvDip ([151]) for EEG data show that ANNs have the potential to yield innovative solutions to the complex inverse problem for MEG data.
Methodology-focused papers typically present innovative architectures or visualization tools optimized for MEG analysis. For instance, in ([125]), the authors propose a novel approach utilizing contrastive learning to develop a foundational model for aligning neural data and visual stimuli across multimodal brain activity representations. Additionally, in ([121]). The authors introduce two new classes of foundational models trained on MEG forecasting tasks, achieving state-of-the-art performance on decoding tasks. Note that methodology-focused studies may overlap with decoding-subcategorized studies but are distinguished by their broader focus on introducing innovative techniques to enhance decoding task efficiency rather than addressing a specific decoding problem.
Preprocessing-focused papers present innovative methods to augment traditional MEG analysis pipelines, targeting improvements in signal-to-noise ratio (SNR), artifact removal, and overall data quality. This subcategory includes a pioneering 2016 preprocessing study proposing Time Contrastive Learning (TCL), an unsupervised representation learning technique, as an alternative to classical Independent Component Analysis (ICA) for artifact removal ([109]). Another study [110] presents a highly accurate artifact classifier, leveraging a hybrid architecture combining deep 1-D and 2-D Convolutional Neural Networks (CNNs) to automate the detection and removal of pervasive whole-brain artifacts, including eye blinks, saccades, and cardiac artifacts. Furthermore, the authors in ([111]) propose a Denoising Autoencoder (DAE) to enhance MEG SNR, providing valuable neurophysiological insights while also offering a user-friendly tool for MEG data visualization and preprocessing.
Unfortunately, only nine studies out of the 34 papers surveyed in this category have their code publicly available or available upon request.
3.4.1 The data
Across the papers we reviewed in this category, the number of trials varied from 294 to 620256, and the number of subjects ranged from three to 676. Nine studies did not mention the amount of data used to train their networks, and two did not mention the number of participants. The studies mostly use MEG data as inputs, occasionally incorporating independent components (IC) of the signal, and two studies use connectivity graphs. Most studies focused on solving the source localization problem use anatomical MRI data in combination with the MEG data.
When MEG data is used, it is generally down-sampled to around 250Hz with some exceptions keeping the sampling rate high (2034Hz in [113], and 810Hz in [132]).
Out of the 34 studies classified as "Other", four used data augmentation ([113, 110, 107, 119]) and seven studies involved simulated data ([131, 130, 129, 134, 111, 133, 41]).
Interestingly, papers classified as methods studies only used MEG data to evaluate the performance of specific methods. It is key in such studies to evaluate the generalization potential of the proposed methods by testing them on different kinds of MEG data recordings (task, rest, etc) or MEG data from different datasets. Yet, we found that only nine out of 16 studies tested their method on more than one dataset or data type. ([117, 113, 115, 119, 121, 124, 123, 126, 125]).
3.4.2 The networks
Most Method and preprocessing studies use CNN-based architectures, while most Source localization-oriented studies use recurrent architectures (mostly LSTM). CNN-based architectures include simple CNNs ([105, 110, 107, 104, 108, 106, 116, 117, 114, 129, 120, 123, 124, 146, 125, 126]) but also Attention-Augmented CNNs ([113, 135]) and recurrent CNNs ([115]). The other studies use MLP ([105, 127, 129]), auto-encoders ([111]), CycleGAN ([119], ResNet ([133], or novel architectures such as Multiple Penalized State Space (MPSS) models ([134]).
Validation techniques are critical when it comes to evaluating the performance of novel methods. Depending on the dataset size, cross-validation might be a mandatory step to accurately assess a model’s prediction and generalization capabilities. All but one of the studies in this category mentioned some kind of validation technique. Source localization studies generally use a simple train/valid split that is sometimes bootstrapped. Method and preprocessing studies use a wider variety of validation techniques but can be condensed into either K-Fold or LOSO and simple train/valid/test splits.
Furthermore, we found that thirteen studies compared the performance of their methods to those of other simpler or more common alternatives ( [104, 105, 132, 129, 127, 131, 130, 122, 121, 123, 124, 135, 146]). Source localization studies generally compare the performance of their method to how Minimum Norm Estimate (MNE) or Beamformer would perform on the same data. Recent source localization studies also compare their results to e- or s-LORETA and ConvDip [151].
In this category, interpreting or visualizing the latent space of the networks might be less crucial for tasks like source localization or preprocessing. However, it remains vital in the methods subcategory, primarily when the study aims to introduce a new architectural approach for decoding tasks. As a matter of fact, in the method subcategory, nine out of 16 studies included some visualization of the latent space of their networks. All of them use a different technique: filter visualization ([113]), activation maps ([115]), channel and time contribution maps ([117]), feature visualization ([116]). In the preprocessing subcategory, five out of nine studies visualize at least partly the latent space of their network mainly through the use of grad-CAM ([108, 105, 110]). Interestingly, none of the source localization studies use any interpretation tool.
@@ 4.1 Added value of combining ANN with MEG
ANNs are increasingly used alongside neuroimaging modalities such as fMRI, EEG, and MEG to enhance our understanding of brain function. fMRI is highly valued for its spatial resolution and ability to map cognitive states through blood flow changes. However, its slower temporal resolution limits its capacity to capture rapid neural dynamics. In contrast, both MEG and EEG offer the temporal precision necessary for analyzing fast brain activities. However, MEG distinguishes itself with superior spatial resolution and reduced susceptibility to scalp and skull distortions. Preprocessing steps, such as artifact removal and feature extraction, are crucial for ensuring that the high-dimensional MEG signals are effectively utilized by ANNs. These steps often form the foundation for downstream tasks by improving signal clarity and optimizing data input for architectures like CNNs, RNNs, and Transformers. In addition to preprocessing, ANNs are widely applied in MEG for classification tasks, including decoding, BCIs, clinical diagnostics, and event detection. By leveraging MEG’s temporal precision, ANNs enable detailed analyses of neural synchronization patterns, facilitating insights into cognitive processes such as perception, attention, language processing, and decision-making. For example, in neural classification, MEG’s precise temporal information supports real-time detection of complex states; in BCI applications, it enhances user control through responsive spatiotemporal decoding; in clinical diagnostics, it enables more accurate seizure detection by addressing signal distortions; and in event detection, it allows subtle oscillatory changes to be captured with higher fidelity. Beyond these applications, MEG–ANN integration plays a critical role in modeling studies, where the goal is to explore how MEG signals align with representations learned by neural networks. These studies compare MEG data with internal ANN activations, providing insights into how cortical activity corresponds to hierarchical network processing. MEG’s high temporal and spatial resolution provides a unique advantage in tracking dynamic neural activity, making it particularly well-suited for investigating how the temporal dynamics of ANN activations align with neural processes in the human brain, especially in domains like vision and audition, where precise timing is critical.
@@ 4.2 Current trends and dominant applications of ANNs in MEG
Taken together, the reviewed body of literature reveals a notable surge in publications and highlights the diverse applications of ANNs in MEG data analysis. This reflects not only a growing appreciation for data-driven approaches in neuroscience but also an increased interest in the unique contributions of ANNs to the field. Based on the papers we reviewed, it seems that the excitement about the potential of ANNs for advancing MEG research is distinct from the recent surge in using standard ML techniques. Our review shows that the power and versatility of deep learning is opening up frameworks for exploring MEG data that go beyond advanced statistical analyses. While many of the papers reviewed do indeed leverage the power of ANNs for classification tasks, many studies benefit from other strengths of ANNs. In particular, the reviewed literature shows that modeling studies that compare representations across ANNs and human MEG data seem to be gaining momentum. Methodological developments for MEG data analytics are also increasingly drawing on the strength and versatility of ANNs. Methodological developments for MEG data analytics are increasingly leveraging the strength and versatility of ANNs. This includes the introduction of novel architectures, enhanced data preprocessing techniques such as artifact detection and correction, innovative source reconstruction approaches, and the development of foundational models. These methodological innovations not only broaden the scope of MEG applications but also accelerate the pace of MEG research. That said, this review also highlights that there are still several critical limitations, including issues with interpretability and data scarcity, which will need to be addressed for the field to harness the potential of ANNs in MEG analysis fully. Further advancements - possibly including more robust foundation models for MEG - are necessary to improve model transparency and effectively manage limited datasets to achieve more reliable and generalizable results. In the next section, we will explore in greater detail the bottlenecks and challenges that emerge from the corpus of studies included in this review.

### v3

@@ 3.4 Other studies
3.4.1 Study aims and subcategories
The final group of studies reviewed includes all MEG-related research involving artificial neural networks that do not fall neatly into either the ’Classification’ or ’Modeling’ categories. These works span a broad range of objectives and methodologies, including preprocessing pipelines, source localization, and the development of novel methods or architectures. Despite their diversity, these studies share a common goal: improving the utility, interpretability, and methodological foundations of MEG analysis using artificial neural networks. This category includes 33 studies and can be divided into three broad subcategories.
The first subcategory focuses on preprocessing techniques [144, 143, 146, 166, 142, 141, 147, 148], proposing ANN-based tools to enhance signal quality by improving artifact detection, SNR, or visualization of neural dynamics. These studies often frame their contributions as supplements to or replacements for traditional techniques like ICA or wavelet filtering.
The second subcategory encompasses studies addressing the source localization problem, also known as the MEG inverse problem [134, 167, 136, 138, 135, 168, 139, 137, 140]. These studies typically train ANNs, often using simulated MEG data, to predict source locations and commonly evaluate performance by comparison to established inverse modeling techniques such as Minimum Norm Estimate (MNE) [195], Beamforming [196], or sLORETA [197].
The final group includes methods-oriented studies [156, 151, 149, 150, 152, 155, 153, 160, 154, 162, 158, 159, 161, 164, 163, 157] that introduce new architectures, training strategies, or analysis frameworks designed to advance ANN-based MEG research. These studies often overlap with classification tasks but are primarily methodological in focus, aiming to improve neural representation learning, domain adaptation, or multimodal alignment rather than solving specific neuroscientific questions.
Although the aims of the studies in this category are heterogeneous, they collectively demonstrate the potential of ANNs to enhance the entire MEG analysis pipeline—from raw data handling to high-level inference. In the following sections, we outline the typical processing workflows used across these studies and examine their data, architectures, training strategies, and evaluation frameworks.
3.4.2 The pipeline
Given the diverse aims of the studies in this category, the pipelines vary considerably depending on whether the study focuses on preprocessing, source localization, or methodological innovation. Nonetheless, each subcategory follows a relatively coherent structure that reflects its specific objectives. Figure 4 (right panel) illustrates a generalized pipeline applicable to many ’Other’-categorized studies.
In preprocessing-focused studies, the pipeline generally begins with raw MEG data acquisition followed by segmentation and optional band-pass filtering. These signals are then passed to an ANN architecture trained to detect and correct artifacts, enhance signal quality, or extract meaningful representations. For instance, some studies used deep convolutional networks to classify time-series segments as clean or contaminated by artifacts such as blinks, saccades, or heartbeats [147], while others proposed autoencoder-based approaches to denoise the data and simultaneously improve interpretability [148].
In source localization studies, the pipeline often begins with simulated dipolar sources projected to sensor space using forward models. These simulated signals, sometimes mixed with noise at different levels, serve as input for training ANNs tasked with predicting the original cortical source locations. Once trained, these models are validated either on additional simulated datasets or real MEG recordings. The ANN output typically includes spatial maps or coordinate predictions that are compared against known ground truths or the output of classical methods like MNE, Beamformers, or sLORETA [134, 167, 136, 138, 135, 168, 139, 137, 140].
Methods-focused studies typically design and validate new architectural or learning frameworks, often repurposing existing MEG datasets for benchmarking. The pipeline usually involves adapting a neural network for a specific task—such as decoding, temporal forecasting, or multimodal alignment—and then comparing its performance to standard models or techniques. In these studies, the output may be a classification, a reconstructed signal, a latent representation, or a learned alignment between modalities [156, 151, 149, 150, 152, 155, 153, 160, 154, 162, 158, 159, 161, 164, 163, 157]. In some cases, these architectures are trained to generalize across datasets or tasks, with evaluation metrics designed to test their robustness, transferability, or explanatory power.
Despite the variability in input-output goals across these pipelines, what unites them is the emphasis on enhancing MEG data processing through neural network-driven components. Whether applied to raw data, source estimation, or methodological refinement, ANNs are increasingly used as flexible tools capable of improving the accuracy, interpretability, or automation of MEG workflows.
3.4.3 The data and preprocessing
The types of data used in this category are as varied as the study objectives, encompassing raw MEG recordings, independent components (ICs), connectivity graphs, and even anatomical MRI when required for source modeling. Most preprocessing and method studies worked directly with raw or minimally processed MEG signals, while source localization studies typically relied on simulated data for training and anatomical priors for real-world testing.
Across the 33 studies in this group, the number of trials used ranged widely, from as few as 294 to over 620,000, and the number of subjects varied between three and 676. However, nine studies did not report the size of their dataset, and two omitted the number of participants. When MEG data were used, downsampling was common. In most cases, the data were resampled to around 250 Hz (see figure 3, although some studies retained higher sampling rates—such as 2000 Hz [140] or 2034 Hz [151]—to preserve fine-grained temporal information. Preprocessing routines generally included band-pass filtering and artifact removal, although detailed procedures were not always specified.
In preprocessing-oriented studies, the raw signals were typically cleaned using either manual ICA, automatic artifact detection algorithms, or ANN-based classifiers. For example, one study introduced Time Contrastive Learning (TCL) as an unsupervised alternative to traditional ICA for separating neural from non-neural sources [141], while another employed a hybrid deep learning architecture combining 1D and 2D CNNs to detect and remove a wide range of artifacts, including eye and cardiac activity [147]. Others used denoising autoencoders to enhance SNR and support visualization of the signal structure [148].
In the source localization subcategory, seven studies explicitly used simulated datasets for training, leveraging ground truth source positions to supervise the learning process [138, 136, 167, 168, 148, 137, 73]. These datasets were often augmented with varying levels of noise to ensure robustness. After training, the models were validated on real MEG recordings, with or without coregistered MRI data. Anatomical information was used to refine spatial accuracy, either by constraining predictions to the cortical surface or by integrating MRI-based head models into the forward projection process.
While some preprocessing and methods papers reused data from previous decoding studies, others evaluated their approach across multiple datasets to test generalizability. Despite this opportunity, only nine of the 16 methods-focused studies applied their model to more than one dataset or task type [155, 151, 150, 160, 162, 161, 159, 163, 164]. This highlights the ongoing challenge of establishing robust benchmarks in ANN-based MEG methodology research.
3.4.4 Network architectures
The diversity of goals in this category is reflected in the wide variety of network architectures employed. While CNNs dominated preprocessing and methodological studies, source localization studies tended to rely on architectures with temporal modeling capabilities such as recurrent networks.
In preprocessing and methods papers, CNN-based architectures were the most commonly used. These included both standard 1D or 2D CNNs [143, 147, 166, 144, 142, 146, 152, 155, 149, 167, 154, 159, 161, 140, 164, 163] and more specialized variants such as attention-augmented CNNs [151, 139] or recurrent CNNs [150]. The choice of CNNs is often motivated by their ability to capture spatiotemporal patterns in MEG signals, and in some cases, to operate directly on time-frequency representations or connectivity matrices. Autoencoders were also employed, especially in unsupervised preprocessing pipelines aiming to denoise or reconstruct input signals [148].
In source localization studies, architectures that explicitly modeled temporal dynamics were more common. These included Long Short-Term Memory (LSTM) networks and other types of Recurrent Neural Networks (RNNs), which are well-suited to tracking changes in MEG activity over time. Some studies also explored ResNet-based models [137] or more novel frameworks such as the multiple penalized state space (MPSS) architecture, designed to integrate spatial and temporal priors [168]. A few studies combined recurrent and convolutional elements to capture both local and distributed features in the MEG signals.
Method-focused studies, in contrast, emphasized innovation in architecture or training paradigms. For instance, one study introduced a contrastive learning framework to align neural and visual representations across modalities [164], while others trained foundational models for MEG forecasting and decoding, achieving state-of-the-art performance across several benchmarks [162]. Another novel direction was the use of generative adversarial networks, such as CycleGAN, to enhance signal translation between domains or conditions [160].
Across all subcategories, the architectures were tailored to specific goals—whether improving signal quality, solving inverse problems, or developing transferable decoding systems. However, these studies also highlighted that ANN architecture alone is rarely sufficient: success often depends on the thoughtful combination of network design, data preparation, and evaluation strategies.
3.4.5 Training and validation techniques
Training and validation procedures varied across studies in this category, depending largely on the objective and data type used. Nonetheless, most studies acknowledged the importance of robust evaluation to assess generalization performance, especially in the context of methodological innovation or when working with small datasets.
In preprocessing and method-oriented studies, training typically followed standard supervised or unsupervised learning procedures. When labeled data were available, supervised learning was used with optimization algorithms such as Adam or SGD, often in combination with dropout, batch normalization, and early stopping to prevent overfitting. Several studies also employed data augmentation strategies to improve generalizability, including sensor shuffling, temporal jittering, or synthetic data generation [151, 147, 166, 160]. In studies leveraging unsupervised approaches like autoencoders or contrastive learning, training aimed to minimize reconstruction loss or maximize representation alignment rather than task-specific accuracy.
Validation strategies in these studies included simple train/validation/test splits, K-fold cross-validation, and Leave-One-Subject-Out (LOSO) schemes. LOSO was most commonly used when generalizing across participants was crucial or when datasets were small. K-fold and bootstrap resampling methods were employed to ensure robustness, particularly when evaluating architectural innovations across multiple configurations. In methods papers proposing general-purpose architectures, performance was often tested across different tasks or datasets to demonstrate transferability.
In source localization studies, validation was often carried out in two stages. First, models were trained and evaluated on simulated data with known source locations and varying levels of noise, allowing precise quantification of localization error using spatial accuracy metrics or distance from the ground truth. Second, the models were applied to real MEG data, and their outputs were compared to those produced by traditional methods such as MNE, beamforming, or sLORETA [138, 136, 167, 168, 148, 137, 73].
Overall, validation was a key element in determining the reliability and potential applicability of the proposed networks. However, the breadth of tasks and data types makes direct comparisons across studies challenging. Greater consistency in reporting training durations, regularization strategies, and evaluation metrics would be beneficial for future work seeking to benchmark ANN-based pipelines in MEG analysis.
3.4.6 Comparison to classical approaches
A distinguishing feature of many studies in this category is their explicit comparison between ANN-based methods and more traditional MEG analysis techniques. These comparisons were critical for justifying the adoption of neural networks, particularly in areas where classical approaches remain well-established, such as source localization or artifact correction.
In the source localization subcategory, ANN models were frequently benchmarked against established inverse solutions, including minimum norm estimate (MNE) [195], Beamformer [196], and variations of LORETA (e.g., eLORETA, sLORETA) [197]. Metrics used for comparison included localization error, area under the curve (AUC), and spatial dispersion. Several studies showed that ANN-based models —particularly those trained on simulated data— were capable of achieving localization accuracy that matched or exceeded classical methods, especially in low signal-to-noise regimes or when dealing with complex source configurations [138, 136, 167, 168, 148, 137, 73]. Recent contributions, such as ConvDip [198], demonstrated that convolutional architectures can provide competitive performance while being more robust to noise and variability in source configurations.
In preprocessing-focused studies, the goal was often to improve artifact detection and signal enhancement beyond what traditional ICA or heuristic thresholding could achieve. For instance, one study replaced ICA with time contrastive learning (TCL), an unsupervised deep learning framework that automatically separated neural from non-neural sources [141]. Another employed a deep CNN-based artifact classifier that outperformed manual rejection and conventional statistical approaches in identifying blinks, heartbeats, and other common contaminants [147].
In the methods subcategory, comparisons to classical baselines were essential for demonstrating the added value of architectural or algorithmic innovations. Thirteen studies explicitly compared their ANN models to simpler or more common alternatives [144, 143, 135, 167, 157, 138, 136, 158, 162, 159, 161, 139, 140]. These comparisons often focused on decoding accuracy, robustness to inter-subject variability, and computational efficiency. Some studies emphasized that classical models require extensive feature engineering or hand-tuned pipelines, while neural networks can learn directly from the data and generalize more flexibly across tasks and domains.
Despite the promising results reported in many of these comparisons, a number of studies also highlighted the interpretability and reproducibility challenges inherent to ANN-based methods. As such, the integration of neural networks into the broader MEG analysis landscape is often framed not as a wholesale replacement of classical techniques, but as a complementary approach that may offer advantages under specific conditions.
@@ 4.1 Added value of combining ANN with MEG
ANNs are increasingly used alongside neuroimaging modalities such as fMRI, EEG, and MEG to enhance our understanding of brain function. fMRI is highly valued for its spatial resolution and ability to map cognitive states through blood flow changes ([9]). However, its slower temporal resolution limits its capacity to capture rapid neural dynamics. In contrast, both MEG and EEG offer the temporal precision necessary for analyzing fast brain activities. However, MEG distinguishes itself with superior spatial resolution and reduced susceptibility to scalp and skull distortions ([8, 12, 6]).
Preprocessing steps, such as artifact removal and feature extraction, are crucial for ensuring that the high-dimensional MEG signals are effectively utilized by ANNs. These steps often form the foundation for downstream tasks by improving signal clarity and optimizing data input for architectures like CNNs, RNNs, and transformers ([142, 143, 141]). In addition to preprocessing, ANNs are widely applied in MEG for classification tasks, including decoding ([118, 59]), BCIs ([93, 103]), clinical diagnostics ([104, 111]), and event detection ([110, 51]). By leveraging MEG’s temporal precision, ANNs have been used to analyze neural synchronization patterns, facilitating insights into cognitive processes such as perception, attention, language processing, and decision-making. For example, in neural classification, MEG’s precise temporal information supports real-time detection of complex states ([97, 55]); in BCI applications, it enhances user control through responsive spatiotemporal decoding ([103, 93]); in clinical diagnostics, it enables more accurate seizure detection by addressing signal distortions; and in event detection ([83, 104]), it allows subtle oscillatory changes to be captured with higher fidelity ([110, 151]).
Beyond these applications, MEG–ANN integration plays a critical role in modeling studies, where the goal is to explore how MEG signals align with representations learned by neural networks. These studies compare MEG data with internal ANN activations, providing insights into how cortical activity corresponds to hierarchical network processing ([119, 122, 120]). MEG’s high temporal and spatial resolution provides a unique advantage in tracking dynamic neural activity, making it particularly well-suited for investigating how the temporal dynamics of ANN activations align with neural processes in the human brain, especially in domains like vision and audition, where precise timing is critical ([118, 129]).
@@ 4.2 Current trends and dominant applications of ANNs in MEG
Taken together, the reviewed body of literature reveals a notable surge in publications and highlights the diverse applications of ANNs in MEG data analysis. This reflects not only a growing appreciation for data-driven approaches in neuroscience but also an increased interest in the unique contributions of ANNs to the field. Based on the papers we reviewed, it seems that the excitement about the potential of ANNs for advancing MEG research is distinct from the recent surge in using standard ML techniques ([5]).
Our review shows that the power and versatility of deep learning is opening up frameworks for exploring MEG data that go beyond advanced statistical analyses. While many of the papers reviewed do indeed leverage the power of ANNs for classification tasks ([118, 97, 93]), many studies benefit from other strengths of ANNs. In particular, the reviewed literature shows that modeling studies that compare representations across ANNs and human MEG data seem to be gaining momentum ([120, 128, 130]).
Methodological developments for MEG data analytics are increasingly leveraging the strength and versatility of ANNs, including the introduction of novel architectures, enhanced data preprocessing techniques such as artifact detection and correction ([141, 144]), innovative source reconstruction approaches ([167, 137]), and the development of foundation models ([162, 114, 199]). These methodological innovations not only broaden the scope of MEG applications but also accelerate the pace of MEG research. That said, this review also highlights that there are still several critical limitations, including issues with interpretability and data scarcity, which will need to be addressed for the field to harness the potential of ANNs in MEG analysis fully. Further advancements —possibly including more robust foundation models for MEG that better handle limited datasets and enhance generalizability— are necessary to improve model transparency and effectively manage limited datasets to achieve more reliable and generalizable results. In the next section, we will explore in greater detail the bottlenecks and challenges that emerge from the corpus of studies included in this review.